OpenStudy (anonymous):

What are the coordinates of the midpoint between the points (-5, 2) and (-5, 8)?

OpenStudy (anonymous):

(-5,6)

OpenStudy (anonymous):

Err (-5,5)

OpenStudy (anonymous):

(8-2)/2 = 3. 2+3 = 5

myininaya (myininaya):

midpoint=((-5+(-5))/2,(2+8)/2)=(-10/2,10/2)=(-5,5)